The nucleus of Element X is unstable and decays into Element Y with a probability z. You are given a pure sample of Element X for a physics assignment, but by the time you get around to doing it, there are six times as many atoms of Element Y as Element X. How much time did you wait before observing in terms of z
Answer #1

This can be done very easily using the concept of the radioactive decay probability distribution.

Given XY het N ~ be the initial number of x. n be the number of decayed a nuleuses out of the total N , it is eanal to the nue , the number of Y atoms at a time t = n = NzE so the number og X atoms remaining to decay - N(t) = N-n =N-Nzt Nuct = N{1-2t